# Towards Federated Graph Learning in One-shot Communication

### v6

@@ A.6.3 Critical Analysis of Current Related Work
Existing FGL methods typically follow the conventional paradigm in which the server adaptively aggregates client models. Under practical non-IID scenarios, numerous studies have shown that this paradigm can lead to issues such as weight divergence [65], client drift [29], and biases and conflicts among client models [70]. These challenges necessitate extensive communication rounds to achieve acceptable performance.
@@
Existing OFL methods are not designed for graph data.
Existing OFL methods are mostly designed for image data. When dealing with graph data, these methods are incompatible or impractical. To be specific, ensemble-based methods such as DENSE train a generator to generate pseudo images via model inversion. However, it’s not perfectly compatible with graph data generation since it can hardly capture the fine-grained structural information of nodes. Distillation-based methods generate distilled data on clients, and the server assembles the distilled data. However, the nodes in a graph are not independent. It’s tricky to assemble generated graphs from clients on the server side. Generative-based methods use VAE or pretrained Stable Diffusion, which cannot generate graph data. Thus, there are not OFL methods that fit perfectly for graph data.
@@
Standalone is a non-federated learning method where each client trains its model only use its local graph data, without communication.
FedAvg is the fundamental federated learning method where each client trains its local model and upload to the server. The server weighted average the client models’ parameters to obtain the global model and then distributes the global model to clients. In our experiments, we limit the upload-download communication round to one round. For better performance, we perform additional finetuning on local graph data after downloading the global model.
FedPUB is a personalized federated graph learning method. FedPUB computes similarities between local models via functional embeddings derived from random graphs generated in the server. The similarities are used for weighted average in personalized aggregation. Additionally, FedPUB employs a personalized mask for each client model to selectively upload parameters. In our experiments, we limit the upload-download communication round to one round.
FedGTA is a personalized federated graph learning method. FedGTA compute and upload the local smoothing confidence and mixed moments of neighbor features to the server to facilitate the weighted average in personalized aggregation. In our experiments, we limit the upload-download communication round to one round.
FedTAD computes the class-wise knowledge reliability scores on each client. The scores, along with local models, are uploaded to the server for data-free knowledge distillation from the ensemble of local models to the global model or personalized model. In our experiments, we limit the upload-download communication round to one round. For better performance, we perform additional finetuning on local graph data after downloading the model.
FedSpray focuses on the minor classes. It optimizes the structure proxy along with the local model on each client. FedSpray uploads and aggregate the structure proxy, which is used in local model training to prevent overfitting and ensure the generalization to the minor classes. In our experiments, we limit the upload-download communication round to one round.
@@
DENSE is a one-shot federated learning method. Each client uploads its trained local model to the server. The server ensembles the client models to generate pseudo data via model inversion, which are used to distill knowledge from client models to the global model or personalized models. For better performance, we perform additional finetuning on local graph data after downloading the data. Note that it’s designed for the image classification task. In our node classification task on graphs, the generators in model inversion on the server-side are trained to generate node features, and the topology structure is constructed using the -Nearest Neighbors strategy as outlined in [69].
Co-Boost is a one-shot federated learning method. It advances the DENSE. The server ensembles the client models with learnable weights. The ensemble weights and generated pseudo data are optimized alternately. For better performance, we perform additional finetuning on local graph data after downloading the data. Note that it’s designed for the image classification task. In our node classification task on graphs, the generators in model inversion on the server-side are trained to generate node features, and the topology structure is constructed using the -Nearest Neighbors strategy as outlined in [69].
A.8 Our method’s Advantages in One Communication Round
@@
A.9 Experimental Results on a Large Number of Clients
@@
A.10 Experimental Results on Link Prediction Task

### v7

@@ H.3 Critical Analysis of Current Related Work
Existing FGL methods typically follow the conventional paradigm in which the server adaptively aggregates client models. Under practical non-IID scenarios, numerous studies have shown that this paradigm can lead to issues such as weight divergence (Zhao et al. 2018), client drift (Karimireddy et al. 2020), and biases and conflicts among client models (Zhu, Hong, and Zhou 2021). These challenges necessitate extensive communication rounds to achieve acceptable performance.
@@
FedGM and GHOST offer methods for training a generalized model in one-shot communication. However, they are not robust to non-IID scenarios. FedGM adopts gradient matching on each client, which is unstable under biased and heterogeneous local graph data. GHOST constructs the pseudo graph with the aligned proxy model, but the proxy model is still biased and heterogeneous under non-IID client graph data. And the performance degrades easily in our experiments. Moreover, neither of them is scalable with respect to the time complexity and memory consumption. The high complexity hinders them to adapt to larger graphs effectively.
Most OFL methods are not designed for graph data.
Existing OFL methods are mostly designed for image data. When dealing with graph data, these methods are incompatible or impractical. To be specific, ensemble-based methods such as DENSE train a generator to generate pseudo images via model inversion. However, it’s not perfectly compatible with graph data generation since it can hardly capture the fine-grained structural information of nodes. Distillation-based methods generate distilled data on clients, and the server assembles the distilled data. However, the nodes in a graph are not independent. It’s tricky to assemble generated graphs from clients on the server side. Generative-based methods use VAE or pretrained Stable Diffusion, which cannot generate graph data. Thus, there are no OFL methods that fit perfectly for graph data.
@@
Standalone is a non-federated learning method where each client trains its model only using its local graph data, without communication.
FedAvg is the fundamental federated learning method where each client trains its local model and uploads it to the server. The server weights averages the client models’ parameters to obtain the global model and then distributes the global model to clients. In our experiments, we limit the upload-download communication round to one round. For better performance, we perform additional fine-tuning on local graph data after downloading the global model.
FedPUB is a personalized federated graph learning method. FedPUB computes similarities between local models via functional embeddings derived from random graphs generated on the server. The similarities are used for weighted average in personalized aggregation. Additionally, FedPUB employs a personalized mask for each client model to selectively upload parameters. In our experiments, we limit the upload-download communication round to one round.
FedGTA is a personalized federated graph learning method. FedGTA computes and uploads the local smoothing confidence and mixed moments of neighbor features to the server to facilitate the weighted average in personalized aggregation. In our experiments, we limit the upload-download communication round to one round.
FedTAD computes the class-wise knowledge reliability scores on each client. The scores, along with local models, are uploaded to the server for data-free knowledge distillation from the ensemble of local models to the global model or personalized model. In our experiments, we limit the upload-download communication round to one round. For better performance, we perform additional fine-tuning on local graph data after downloading the model.
FedSpray focuses on the minor classes. It optimizes the structure proxy along with the local model on each client. FedSpray uploads and aggregates the structure proxy, which is used in local model training to prevent overfitting and ensure the generalization to the minor classes. In our experiments, we limit the upload-download communication round to one round.
@@
FedGM adopts gradient matching to condense the synthesized graph on each client. Each client uploads the synthesized graphs or gradients to the server, and the server aggregates them. In our experiment, we follow the setting from the original paper to conduct one-round communication and use aggregated synthesized graphs to aid the model training. For better performance, we perform additional fine-tuning on local graph data after downloading the global model.
GHOST aims to train a generalized model in one-shot communication. Each client constructs a pseudo graph and trains a proxy model by aligning the pseudo graph and the real local graph. Each client then uploads the proxy model. The server generates the pseudo-graph with each uploaded proxy model and trains a global model with crucial parameters stabilized to alleviate catastrophic forgetting. For better performance, we perform additional fine-tuning on local graph data after downloading the global model.
DENSE is a one-shot federated learning method. Each client uploads its trained local model to the server. The server ensembles the client models to generate pseudo data via model inversion, which is used to distill knowledge from client models to the global model or personalized models. For better performance, we perform additional fine-tuning on local graph data after downloading the data. Note that it’s designed for the image classification task. In our node classification task on graphs, the generators in model inversion on the server-side are trained to generate node features, and the topology structure is constructed using the -Nearest Neighbors strategy as outlined in (Zhu et al. 2024b).
Co-Boost is a one-shot federated learning method. It advances the DENSE. The server ensembles the client models with learnable weights. The ensemble weights and generated pseudo data are optimized alternately. For better performance, we perform additional fine-tuning on local graph data after downloading the data. Note that it’s designed for the image classification task. In our node classification task on graphs, the generators in model inversion on the server-side are trained to generate node features, and the topology structure is constructed using the -Nearest Neighbors strategy as outlined in (Zhu et al. 2024b).
Appendix J Our method’s Advantages in One Communication Round
@@
Appendix K Experimental Results on a Large Number of Clients
@@
Appendix L Experimental Results on Link Prediction Task